# Local Advice and Local Decompression

### v1

@@ 6.2. $\Delta$-coloring
Algorithm: Our algorithm can first recolor all vertices in local neighborhoods of relay vertices. To that end, all vertices explore their $O(\alpha\log\Delta)$ neighborhoods and reconstruct the shortest path trees around the relay nodes. Then, the algorithm uses the same technique as in (Panconesi and Srinivasan, 1992) (process vertices on paths, leaf to root), but instead of a randomized step that breaks the symmetry, it uses Theorem 8 as in the algorithm from (Ghaffari et al., 2021). In total, this step is the most time-consuming part of the algorithm and takes $O(\alpha\log\Delta\cdot\sqrt{\Delta\log\Delta})$ time. The result is that all non-relay nodes that are not in $R$ were colored.
Then, the relay vertices can assign to themselves a new color (decoded from the advice bits scattered in its $O(\alpha\log\Delta)$ hop neighborhood), which leaves us with the task of re-coloring the vertices on the paths between the relay nodes. Each relay node knows its ancestor in the previous layer of relays, as such the vertices on the paths that need recoloring can be identified, and new colors can be assigned. ∎
Fixing coloring of root vertices
Finally, let us consider the last subproblem, of extending partial $\Delta$-coloring into proper $\Delta$-coloring, given that the uncolored vertices are at a distance larger than $\Omega(\log_{\Delta}n)$. The standard algorithms use a somewhat brute-force approach, that for each uncolored vertex finds a sequence of recolorings that if applied, results in a proper $\Delta$-coloring. By Lemma 7, a set of such nonoverlapping sequences can be found by exploring $O(\log_{\Delta}n)$ hop neighborhoods of all uncolored vertices. Once again, one can see this sequence of recolorings as a sequence of dependent problems and the advice can be used to both define the sequence and make the dependency chains shorter.
Problem 4.
We are given a graph $G$ with partial $\Delta$-coloring $C$, such that any two vertices $u,v$ such that $C(v)=\bot$ and $C(u)=\bot$ are at distance larger than $2c\log_{\Delta}n$, where $c$ is a constant from Lemma 7. The goal is to compute a proper $\Delta$-coloring of $G$.
Lemma 0.
Let $\mathcal{G}_{\Delta}$ be the set of all $\Delta$-colorable graphs with vertices partially colored according to some proper $\Delta$-coloring, in which any two uncolored vertices are at a distance larger than $2c\log_{\Delta}n$. Let $\Pi_{\Delta}$ be the problem described in 3. Then, there exists a $(\mathcal{G}_{\Delta},\Pi_{\Delta},\gamma_{0},A,T)$-composable advice schema, where $\gamma_{0}=O(1)$, $A(c,\gamma)=\Theta(\gamma^{3})$, and $T(\alpha,\Delta)=O(\alpha)$.
Proof.
@@
By Lemma 7 we know that there exists a set of vertices $X$ such that each uncolored vertex in $G$ has a vertex from $X$ such that $G$ can be properly colored by shifting the colors along the shortest paths from uncolored vertices to vertices from $X$ and recoloring vertices in $X$ with a free color. Our composable advice schema puts:
@@
  • markers $11$ on every $102\alpha+$ 10th vertex on the shortest paths from the uncolored vertices to vertices from $X$ (leaving one block of unmarked vertices of length slightly longer, but at most two times longer, to accommodate for the fact that the length of the path may be not divisible by $102\alpha+$ 10),
@@
Decoding algorithm: to compute the paths, each vertex with no color or with marker $11$ gathers its $222\alpha+$ 22 hop neighborhood. For each vertex, such information is sufficient to identify $222\alpha+$ 22 predecessors and successors on the shortest path between uncolored vertices and $X$, and a direction towards $X$. This information can be then broadcasted by each vertex marked by $11$ to all vertices on the path. Then, the path vertices can update their color accordingly.
As for vertices with marker $111$, they wait for other vertices to recolor and then recolor themselves. The property of vertices was that either they have a low degree (and then they always can recolor themselves with a color no larger than $\Delta$), or that they have two neighbors of the same color, none being on the selected shortest paths from $I$ to $X$. As such, after recoloring the vertices from $X$ still have at least one free color to choose from (as they still have degrees smaller than $\Delta$ or two neighbors with the same color). ∎
7. 3-Coloring 3-Colorable Graphs
In this section, we devise an algorithm that colors any $3$-colorable graph with $3$ colors in $\operatorname{poly}(\Delta)$ communication rounds, given $1$ bit of advice per node. More in detail, we prove the following theorem.
Theorem 1.
@@
Proof.
Assume there is no node $w$ with the stated property (otherwise we are done). This in particular implies that any node in distance at most $\Delta$ from $v$ in $C$ has exactly one neighbor of color $1$ in $G$ (as the given $3$-coloring is a greedy $3$-coloring). For $v$, let $u$ denote this unique neighbor. Consider any path $v=v_{0},v_{1},\dots,v_{\Delta}$ of length $\Delta$ starting in $v$. (Such a path must exist as otherwise the diameter of $C$ would be strictly smaller than $2\Delta$.) As $u$ has degree at most $\Delta$ and $v$ is a neighbor of $u$, at least one node from $\{v_{1},\dots,v_{\Delta}\}$ is not a neighbor of $u$. This implies that there is some index $10\leq i\leq\Delta-$ 1 such that exactly one of $v_{i}$ and $v_{i+1}$ is a neighbor of $u$. As both $v_{i}$ and $v_{i+1}$ have exactly one neighbor of color $1$ in $G$, it follows that $v_{i}$ and $v_{i+1}$ constitute nodes $x$ and $y$ with the properties described in the lemma. ∎
Now we are set to formally describe our encoding schema. While it might not be immediately obvious why some objects that the schema computes along the way exist, we will show subsequently in Lemma 3 that they indeed exist.
@@
Fix an arbitrary greedy $3$-coloring $\varphi\colon V(G)\rightarrow\{1,2,3\}$ of the nodes of $G$ (i.e., each node of color $i$ has neighbors of all colors $<i$). Assign bit $1$ to all nodes of color $1$.
@@
Select either a node $w$ or two nodes $x,y$ as described in Lemma 2, applied to $v$. Denote the set of selected nodes by $S_{v}$ (i.e., either $S_{v}=\{w\}$ or $S_{v}=\{x,y\}$). Denote by $T_{v}$ the set of all nodes of $V(C)$ that
  1. (1)
@@
Select either a node $w^{\prime}$ or two nodes $x^{\prime},y^{\prime}$ as described in Lemma 2, applied to $v^{\prime}$, with the additional constraint that the selected node or a pair of nodes must be contained in $P_{v}$. Denote the set of selected nodes by $S^{\prime}_{v}$ (i.e., either $S^{\prime}_{v}=\{w^{\prime}\}$ or $S^{\prime}_{v}=\{x^{\prime},y^{\prime}\}$).
@@
Finally, assign bit 00 to all nodes that have not been assigned bit $1$ so far.
@@
Next, the existence of a node $w$ or two nodes $x,y$ as described follows from Lemma 2.
Now, consider node $v^{\prime}$ and path $P_{v}$. We begin the proof of their existence by bounding the number of nodes in $V(C)\setminus T_{v}$. Observe that there are at most $2\Delta^{2}$ nodes that share a neighbor of color $1$ with some node from $S_{v}$ since $|S_{v}|\leq 2$. Observe further that $|S_{v}|\leq 2$ also implies that there are at most $22\Delta+$ 2 nodes that are neighbors of or identical to some node in $S_{v}$. Hence, $|V(C)\setminus T_{v}|\leq 2\Delta^{2}+2\Delta+2\leq 6\Delta^{2}$. Now consider a node $v_{20\Delta^{3}}$ that is at distance exactly $20\Delta^{3}$ from $v$ in $C$ (whose existence again follows from the fact that $C$ has diameter larger than $4000\Delta^{9}$). Let $(v=v_{0},v_{1},v_{2},\dots,v_{20\Delta^{3}})$ be a shortest path from $v$ to $v_{20\Delta^{3}}$. Since $|V(C)\setminus T_{v}|\leq 6\Delta^{2}$, there must be a subpath $(v_{i},v_{i+1},\dots,v_{i+\Delta})$ of length $\Delta$ such that all nodes on the subpath are contained in $T_{v}$. Therefore, by setting $v^{\prime}:=v_{i}$ and $P_{v}:=(v_{i},v_{i+1},\dots,v_{i+\Delta})$, we ensure that $v^{\prime}$ and $P_{v}$ satisfy the properties stated in the encoding schema, concluding the existence proof for $v^{\prime}$ and $P_{v}$.
To show the existence of a node $w^{\prime}$ or two nodes $x^{\prime},y^{\prime}$ as described in the schema, it suffices to observe that the proof of Lemma 2 also works with the additional restriction that the selected node $w^{\prime}$ or the selected two nodes $x^{\prime},y^{\prime}$ are contained in $P_{v}$ (as the proof, in fact, starts by finding such a path).
@@
Next, we bound the number of random variables an event $A_{u}$ depends on. Observe that for each set $Q_{r}$ (where $r\in R_{C}$ for some $C\in\mathcal{C}(G^{*}_{2,3})$), each node $v\in Q_{r}$, and each node $w\in S_{v}\cup S^{\prime}_{v}$, we have $\operatorname{dist}_{C}(r,w)\leq\operatorname{dist}_{C}(r,v)+\operatorname{dist}_{C}(v,w)\leq 600\Delta^{9}+20\Delta^{3}+\Delta\leq 700\Delta^{9}$ (by Lemma 2 and the distance properties stated in the description of the encoding schema). Hence, for any two different pairs $(C,r)\neq(C^{\prime},r^{\prime})$ (as specified above), and any two nodes $u\in Q_{r}$, $v\in Q_{r^{\prime}}$, we have $(S_{u}\cup S^{\prime}_{u})\cap(S_{v}\cup S^{\prime}_{v})=\emptyset$: either we have $C\neq C^{\prime}$, in which case $S_{u}\cup S^{\prime}_{u}\in V(C)$ and $S_{v}\cup S^{\prime}_{v}\in V(C^{\prime})$ are subsets of the node sets of different components, or $C=C^{\prime}$ and $r\neq r^{\prime}$, in which case the computed distance upper bound of $700\Delta^{9}$, together with the fact that the nodes in $R_{C}$ have a pairwise distance of at least $2000\Delta^{9}$ in $C$, ensures that any node in $(S_{u}\cup S^{\prime}_{u})$ has a distance (in $C$) of at least $600\Delta^{9}$ from any node in $(S_{v}\cup S^{\prime}_{v})$. By the definition of the $A_{u}$, it follows that each event $A_{u}$ depends on at most $\Delta$ random variables (as $u$ has at most $\Delta$ neighbors).
@@
Now we are ready to apply the Lovász Local Lemma. By the bounds established above, we know that $epd\leq e\cdot 1/(144\Delta^{8})\cdot 48\Delta^{8}<1$. By the Lovász Local Lemma, this implies that there exists an assignment to the random variables $X_{r,C}$ such that none of the events $A_{u}$ occurs. In other words, it is possible to select, for each $C\in\mathcal{C}(G^{*}_{2,3})$, and each $r\in R_{C}$, a node $v_{r,C}$ such that for each node $u$ of color $1$ there exists at most one pair $(r,C)$ such that $u$ has a neighbor in $(S_{v_{r,C}}\cup S^{\prime}_{v_{r,C}})$. Moreover, for any fixed pair $(r,C)$ and any node $u$ of color $1$, at most one neighbor of $u$ is contained in $(S_{v_{r,C}}\cup S^{\prime}_{v_{r,C}})$, by the guarantees provided by Lemma 2 and the construction of $S^{\prime}_{v}$ (which ensures that $S^{\prime}_{v}\subseteq T_{v}$). Hence, the aforementioned selection of nodes $v_{r,C}$ indeed satisfies the property specified in the encoding schema, concluding the proof. ∎
@@
By the definitions of $S_{v_{r,C}}$ and $S^{\prime}_{v_{r,C}}$ (and the guarantees specified in Lemma 2) and the choice of exactly which nodes of color $\neq 1$ receive a $1$-bit, this implies that $v$ has at least two neighbors of color $1$ or $v$ has a neighbor $u$ of color $\neq 1$ that also receives bit $1$. As we are done in the former case, assume the latter, i.e., that $v$ has a neighbor $u$ with $\varphi(u)\neq 1$ and $b(u)=1$. Since $\varphi$ is a greedy coloring and $\varphi(u)\neq 1$, we know that $v$ has a neighbor $w$ of color $1$ (which therefore satisfies $b(w)=1$). As $\varphi(u)\neq 1$, we have $w\neq u$, which implies that $v$ has two neighbors that receive bit $1$. ∎
@@
Consider first nodes of color $1$ in $\psi$. These nodes have color $1$ also in $\varphi$, and by Lemma 4 and the definition of $\mathcal{A}$, it follows that these nodes output color $1$, as desired.
@@
Finally, let $u$ be a node with $\psi(u)\neq 1$ for which the diameter of $C(u)$ is strictly larger than $4000\Delta^{9}$. Observe that each node in $C(u)$ receives the same color under $\psi$ as under $\varphi$. Consider the node $w$ selected by $u$ in $\mathcal{A}$. By the definition of the encoding schema, there exist nodes $r\in R_{C(u)}$ and $v_{r,C(u)}\in Q_{r}$ such that $w\in(S_{v_{r,C}}\cup S^{\prime}_{v_{r,C}})$. As essentially already observed in the proof of Lemma 3, we know that for any $r^{\prime}\in R_{C(u)}$ satisfying $r^{\prime}\neq r$, and any node $y\in(S_{v_{r^{\prime},C}}\cup S^{\prime}_{v_{r^{\prime},C}})$, we have $\operatorname{dist}_{C(u)}(w,y)\geq 600\Delta^{9}$. By the specification of which nodes receive bit $1$ in the encoding schema, it follows that any node in the $(30\Delta^{3})$-hop neighborhood of $w$ in $C(u)$ that receives bit $1$ is contained in $S_{v_{r,C}}\cup S^{\prime}_{v_{r,C}}$. Vice versa, since any two nodes in $S_{v_{r,C}}\cup S^{\prime}_{v_{r,C}}$ have distance at most $\Delta+20\Delta^{3}+\Delta$ in $C(u)$ and $w\in(S_{v_{r,C}}\cup S^{\prime}_{v_{r,C}})$, all nodes from $S_{v_{r,C}}\cup S^{\prime}_{v_{r,C}}$ are contained in the $(30\Delta^{3})$-hop neighborhood of $w$ in $C(u)$. Hence, it suffices to restrict attention to which nodes from $S_{v_{r,C}}\cup S^{\prime}_{v_{r,C}}$ receive bit $1$ in the encoding schema in order to determine which color $u$ outputs according to $\mathcal{A}$.

### v2

@@ C.2 $\Delta$-coloring
Algorithm: Our algorithm can first recolor all vertices in local neighborhoods of relay vertices. To that end, all vertices explore their $O(\alpha\log\Delta)$ neighborhoods and reconstruct the shortest path trees around the relay nodes. Then, the algorithm uses the same technique as in [53] (process vertices on paths, leaf to root), but instead of a randomized step that breaks the symmetry, it uses Theorem˜C.12 as in the algorithm from [31]. In total, this step is the most time-consuming part of the algorithm and takes $O(\alpha\log\Delta\cdot\sqrt{\Delta\log\Delta})$ time. The result is that all non-relay nodes that are not in $R$ were colored.
Then, the relay vertices can assign to themselves a new color (decoded from the advice bits scattered in its $O(\alpha\log\Delta)$ hop neighborhood), which leaves us with the task of re-coloring the vertices on the paths between the relay nodes. Each relay node knows its ancestor in the previous layer of relays, as such the vertices on the paths that need recoloring can be identified, and new colors can be assigned.
Fixing coloring of root vertices.
Finally, let us consider the last subproblem, of extending partial $\Delta$-coloring into proper $\Delta$-coloring, given that the uncolored vertices are at a distance larger than $\Omega(\log_{\Delta}n)$. The standard algorithms use a somewhat brute-force approach, that for each uncolored vertex finds a sequence of recolorings that if applied, results in a proper $\Delta$-coloring. By Lemma˜C.11, a set of such nonoverlapping sequences can be found by exploring $O(\log_{\Delta}n)$ hop neighborhoods of all uncolored vertices. Once again, one can see this sequence of recolorings as a sequence of dependent problems and the advice can be used to both define the sequence and make the dependency chains shorter.
Problem C.16.
We are given a graph $G$ with partial $\Delta$-coloring $C$, such that any two vertices $u,v$ such that $C(v)=\bot$ and $C(u)=\bot$ are at distance larger than $2c\log_{\Delta}n$, where $c$ is a constant from Lemma˜C.11. The goal is to compute a proper $\Delta$-coloring of $G$.
Lemma C.17.
Let $\mathcal{G}_{\Delta}$ be the set of all $\Delta$-colorable graphs with vertices partially colored according to some proper $\Delta$-coloring, in which any two uncolored vertices are at a distance larger than $2c\log_{\Delta}n$. Let $\Pi_{\Delta}$ be the problem described in ˜C.13. Then, there exists a $(\mathcal{G}_{\Delta},\Pi_{\Delta},\gamma_{0},A,T)$-composable advice schema, where $\gamma_{0}=O(1)$, $A(c,\gamma)=\Theta(\gamma^{3})$, and $T(\alpha,\Delta)=O(\alpha)$.
Proof C.18.
@@
By Lemma˜C.11 we know that there exists a set of vertices $X$ such that each uncolored vertex in $G$ has a vertex from $X$ such that $G$ can be properly colored by shifting the colors along the shortest paths from uncolored vertices to vertices from $X$ and recoloring vertices in $X$ with a free color. Our composable advice schema puts:
@@
  • markers $11$ on every $2\alpha+10$th vertex on the shortest paths from the uncolored vertices to vertices from $X$ (leaving one block of unmarked vertices of length slightly longer, but at most two times longer, to accommodate for the fact that the length of the path may be not divisible by $2\alpha+10$),
@@
Decoding algorithm: to compute the paths, each vertex with no color or with marker $11$ gathers its $2\alpha+22$ hop neighborhood. For each vertex, such information is sufficient to identify $2\alpha+22$ predecessors and successors on the shortest path between uncolored vertices and $X$, and a direction towards $X$. This information can be then broadcasted by each vertex marked by $11$ to all vertices on the path. Then, the path vertices can update their color accordingly.
As for vertices with marker $111$, they wait for other vertices to recolor and then recolor themselves. The property of vertices was that either they have a low degree (and then they always can recolor themselves with a color no larger than $\Delta$), or that they have two neighbors of the same color, none being on the selected shortest paths from $I$ to $X$. As such, after recoloring the vertices from $X$ still have at least one free color to choose from (as they still have degrees smaller than $\Delta$ or two neighbors with the same color).
Appendix D 3-coloring 3-colorable graphs
In this section, we devise an algorithm that colors any $3$-colorable graph with $3$ colors in $\operatorname{poly}(\Delta)$ communication rounds, given $1$ bit of advice per node. In more detail, we prove the following theorem.
Theorem D.1.
@@
Proof D.3.
Assume there is no node $w$ with the stated property (otherwise we are done). This in particular implies that any node in distance at most $\Delta$ from $v$ in $C$ has exactly one neighbor of color $1$ in $G$ (as the given $3$-coloring is a greedy $3$-coloring). For $v$, let $u$ denote this unique neighbor. Consider any path $v=v_{0},v_{1},\dots,v_{\Delta}$ of length $\Delta$ starting in $v$. (Such a path must exist as otherwise the diameter of $C$ would be strictly smaller than $2\Delta$.) As $u$ has degree at most $\Delta$ and $v$ is a neighbor of $u$, at least one node from $\{v_{1},\dots,v_{\Delta}\}$ is not a neighbor of $u$. This implies that there is some index $0\leq i\leq\Delta-1$ such that exactly one of $v_{i}$ and $v_{i+1}$ is a neighbor of $u$. As both $v_{i}$ and $v_{i+1}$ have exactly one neighbor of color $1$ in $G$, it follows that $v_{i}$ and $v_{i+1}$ constitute nodes $x$ and $y$ with the properties described in the lemma.
Now we are set to formally describe our encoding schema. While it might not be immediately obvious why some objects that the schema computes along the way exist, we will show subsequently in Lemma˜D.4 that they indeed exist.
@@
Fix an arbitrary greedy $3$-coloring $\varphi\colon V(G)\rightarrow\{1,2,3\}$ of the nodes of $G$ (i.e., each node of color $i$ has neighbors of all colors <i<i). Assign bit $1$ to all nodes of color $1$.
@@
Select either a node $w$ or two nodes $x,y$ as described in Lemma˜D.2, applied to $v$. Denote the set of selected nodes by $S_{v}$ (i.e., either $S_{v}=\{w\}$ or $S_{v}=\{x,y\}$). Denote by $T_{v}$ the set of all nodes of $V(C)$ that
  1. 1.
@@
Select either a node $w^{\prime}$ or two nodes $x^{\prime},y^{\prime}$ as described in Lemma˜D.2, applied to $v^{\prime}$, with the additional constraint that the selected node or a pair of nodes must be contained in $P_{v}$. Denote the set of selected nodes by $S^{\prime}_{v}$ (i.e., either $S^{\prime}_{v}=\{w^{\prime}\}$ or $S^{\prime}_{v}=\{x^{\prime},y^{\prime}\}$).
@@
Finally, assign bit 0 to all nodes that have not been assigned bit $1$ so far.
@@
Next, the existence of a node $w$ or two nodes $x,y$ as described follows from Lemma˜D.2.
Now, consider node $v^{\prime}$ and path $P_{v}$. We begin the proof of their existence by bounding the number of nodes in $V(C)\setminus T_{v}$. Observe that there are at most $2\Delta^{2}$ nodes that share a neighbor of color $1$ with some node from $S_{v}$ since $|S_{v}|\leq 2$. Observe further that $|S_{v}|\leq 2$ also implies that there are at most $2\Delta+2$ nodes that are neighbors of or identical to some node in $S_{v}$. Hence, $|V(C)\setminus T_{v}|\leq 2\Delta^{2}+2\Delta+2\leq 6\Delta^{2}$. Now consider a node $v_{20\Delta^{3}}$ that is at distance exactly $20\Delta^{3}$ from $v$ in $C$ (whose existence again follows from the fact that $C$ has diameter larger than $4000\Delta^{9}$). Let $(v=v_{0},v_{1},v_{2},\dots,v_{20\Delta^{3}})$ be a shortest path from $v$ to $v_{20\Delta^{3}}$. Since $|V(C)\setminus T_{v}|\leq 6\Delta^{2}$, there must be a subpath $(v_{i},v_{i+1},\dots,v_{i+\Delta})$ of length $\Delta$ such that all nodes on the subpath are contained in $T_{v}$. Therefore, by setting $v^{\prime}:=v_{i}$ and $P_{v}:=(v_{i},v_{i+1},\dots,v_{i+\Delta})$, we ensure that $v^{\prime}$ and $P_{v}$ satisfy the properties stated in the encoding schema, concluding the existence proof for $v^{\prime}$ and $P_{v}$.
To show the existence of a node $w^{\prime}$ or two nodes $x^{\prime},y^{\prime}$ as described in the schema, it suffices to observe that the proof of Lemma˜D.2 also works with the additional restriction that the selected node $w^{\prime}$ or the selected two nodes $x^{\prime},y^{\prime}$ are contained in $P_{v}$ (as the proof, in fact, starts by finding such a path).
@@
Next, we bound the number of random variables an event $A_{u}$ depends on. Observe that for each set $Q_{r}$ (where $r\in R_{C}$ for some $C\in\mathcal{C}(G^{*}_{2,3})$), each node $v\in Q_{r}$, and each node $w\in S_{v}\cup S^{\prime}_{v}$, we have $\operatorname{dist}_{C}(r,w)\leq\operatorname{dist}_{C}(r,v)+\operatorname{dist}_{C}(v,w)\leq 600\Delta^{9}+20\Delta^{3}+\Delta\leq 700\Delta^{9}$ (by Lemma˜D.2 and the distance properties stated in the description of the encoding schema). Hence, for any two different pairs $(C,r)\neq(C^{\prime},r^{\prime})$ (as specified above), and any two nodes $u\in Q_{r}$, $v\in Q_{r^{\prime}}$, we have $(S_{u}\cup S^{\prime}_{u})\cap(S_{v}\cup S^{\prime}_{v})=\emptyset$: either we have $C\neq C^{\prime}$, in which case $S_{u}\cup S^{\prime}_{u}\in V(C)$ and $S_{v}\cup S^{\prime}_{v}\in V(C^{\prime})$ are subsets of the node sets of different components, or $C=C^{\prime}$ and $r\neq r^{\prime}$, in which case the computed distance upper bound of $700\Delta^{9}$, together with the fact that the nodes in $R_{C}$ have a pairwise distance of at least $2000\Delta^{9}$ in $C$, ensures that any node in $(S_{u}\cup S^{\prime}_{u})$ has a distance (in $C$) of at least $600\Delta^{9}$ from any node in $(S_{v}\cup S^{\prime}_{v})$. By the definition of the $A_{u}$, it follows that each event $A_{u}$ depends on at most $\Delta$ random variables (as $u$ has at most $\Delta$ neighbors).
@@
Now we are ready to apply the Lovász Local Lemma. By the bounds established above, we know that $epd\leq e\cdot 1/(144\Delta^{8})\cdot 48\Delta^{8}<1$. By the Lovász Local Lemma, this implies that there exists an assignment to the random variables $X_{r,C}$ such that none of the events $A_{u}$ occurs. In other words, it is possible to select, for each $C\in\mathcal{C}(G^{*}_{2,3})$, and each $r\in R_{C}$, a node $v_{r,C}$ such that for each node $u$ of color $1$ there exists at most one pair $(r,C)$ such that $u$ has a neighbor in $(S_{v_{r,C}}\cup S^{\prime}_{v_{r,C}})$. Moreover, for any fixed pair $(r,C)$ and any node $u$ of color $1$, at most one neighbor of $u$ is contained in $(S_{v_{r,C}}\cup S^{\prime}_{v_{r,C}})$, by the guarantees provided by Lemma˜D.2 and the construction of $S^{\prime}_{v}$ (which ensures that $S^{\prime}_{v}\subseteq T_{v}$). Hence, the aforementioned selection of nodes $v_{r,C}$ indeed satisfies the property specified in the encoding schema, concluding the proof.
@@
By the definitions of $S_{v_{r,C}}$ and $S^{\prime}_{v_{r,C}}$ (and the guarantees specified in Lemma˜D.2) and the choice of exactly which nodes of color $\neq 1$ receive a $1$-bit, this implies that $v$ has at least two neighbors of color $1$ or $v$ has a neighbor $u$ of color $\neq 1$ that also receives bit $1$. As we are done in the former case, assume the latter, i.e., that $v$ has a neighbor $u$ with $\varphi(u)\neq 1$ and $b(u)=1$. Since $\varphi$ is a greedy coloring and $\varphi(u)\neq 1$, we know that $v$ has a neighbor $w$ of color $1$ (which therefore satisfies $b(w)=1$). As $\varphi(u)\neq 1$, we have $w\neq u$, which implies that $v$ has two neighbors that receive bit $1$.
@@
Consider first nodes of color $1$ in $\psi$. These nodes have color $1$ also in $\varphi$, and by Lemma˜D.6 and the definition of $\mathcal{A}$, it follows that these nodes output color $1$, as desired.
@@
Finally, let $u$ be a node with $\psi(u)\neq 1$ for which the diameter of $C(u)$ is strictly larger than $4000\Delta^{9}$. Observe that each node in $C(u)$ receives the same color under $\psi$ as under $\varphi$. Consider the node $w$ selected by $u$ in $\mathcal{A}$. By the definition of the encoding schema, there exist nodes $r\in R_{C(u)}$ and $v_{r,C(u)}\in Q_{r}$ such that $w\in(S_{v_{r,C}}\cup S^{\prime}_{v_{r,C}})$. As essentially already observed in the proof of Lemma˜D.4, we know that for any $r^{\prime}\in R_{C(u)}$ satisfying $r^{\prime}\neq r$, and any node $y\in(S_{v_{r^{\prime},C}}\cup S^{\prime}_{v_{r^{\prime},C}})$, we have $\operatorname{dist}_{C(u)}(w,y)\geq 600\Delta^{9}$. By the specification of which nodes receive bit $1$ in the encoding schema, it follows that any node in the $(30\Delta^{3})$-hop neighborhood of $w$ in $C(u)$ that receives bit $1$ is contained in $S_{v_{r,C}}\cup S^{\prime}_{v_{r,C}}$. Vice versa, since any two nodes in $S_{v_{r,C}}\cup S^{\prime}_{v_{r,C}}$ have distance at most $\Delta+20\Delta^{3}+\Delta$ in $C(u)$ and $w\in(S_{v_{r,C}}\cup S^{\prime}_{v_{r,C}})$, all nodes from $S_{v_{r,C}}\cup S^{\prime}_{v_{r,C}}$ are contained in the $(30\Delta^{3})$-hop neighborhood of $w$ in $C(u)$. Hence, it suffices to restrict attention to which nodes from $S_{v_{r,C}}\cup S^{\prime}_{v_{r,C}}$ receive bit $1$ in the encoding schema in order to determine which color $u$ outputs according to $\mathcal{A}$.